OpenStudy (anonymous):

Hilda has deposited $874 in a savings account that earns interest at a rate of 3.2% compounded quarterly. What will the account balance be in seven years?

OpenStudy (anonymous):

$901.97 $1,092.46 $2,111.27 $2,796.80

OpenStudy (anonymous):

multiple choice

OpenStudy (johnweldon1993):

\[\large A = P(1 + \frac{ r }{ n })^{nt}\] A = amount (what you want...) P = principal / original rate r = rate (.032 here) n = the number of times the interests gets compounded...."quarterly" means 4 times a year...so 4 and t = time in years... so we have \[\large A = 874(1 + \frac{ .032 }{ 4 })^{4 \times 7}\] Can you solve this?

OpenStudy (anonymous):

Yes i believe i can hold up a second please

OpenStudy (johnweldon1993):

No problem...

OpenStudy (anonymous):

is it 24,667.776

OpenStudy (anonymous):

no i did it wrong right?

OpenStudy (johnweldon1993):

Yeah..show me what you did...

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

i just re calculated

OpenStudy (johnweldon1993):

Yes it is...good work!

OpenStudy (anonymous):

thank you so much

OpenStudy (johnweldon1993):

Anytime!

OpenStudy (anonymous):

bye now (: